OpenStudy (anonymous):

When she turned 21, Reese purchased 125 shares of Stock A for $35 per share; 72 shares of Stock B for $18 per share; 56 shares of Stock C for $43 per share; and 105 shares of Stock D for $26 per share. At age 30, Reese sold Stock A for $4804 with annual dividends of $75 per year, Stock B for $1456, Stock C for $2384, and Stock D for $3475. Which stock provided Reese with the greatest return on her investment? Stock A Stock B Stock C Stock D

OpenStudy (anonymous):

from age 21 to age 30 = 8years Stock a= investment is 125*35 = 4375 proceeds = dividends plus sales = 75*8 + 4804 = 600+4804 = $5404 profit = 5404- 4375 =1029 ..that is 23.52 % on investment. stock b = investment is 72 * 18 = 1296 proceeds = 1456 profit = 1456-1296 = 160, that is 12.35% on investments stock c = investment is 56 * 43 =2408 proceeds = 2384 loss = 2408-2384 = 24 (negative) stock d = investment is 2730 proceeds = 3475 profits = 3475-2730 = 745, that is 27.28 % on investments. it means stock D provided the highest return.